OpenStudy (anonymous):

Mike wants to make meatloaf. His recipe uses a total of 4 pounds of meat. If he uses a 3 to 1 ratio of beef to pork, how much pork will he use? Enter your answer as a mixed number in simplest terms. @agent0smith @mathslover @kelliegirl33 @RICARDOismyfirstname

OpenStudy (anonymous):

3 to 1 ratio means four parts per whole (75% and 25%) So four every pound of meat, he will use a quarter pound of pork and three quarter pounds of beef. .25•6=1.5 and 1.5 is 1 and 1/2 in mixed numbers (Pork) 1.25•6=4.5 and 4.5 is 4 and 1/2 in mixed numbers (Beef) 4.5/1.5=3/1 and 4.5+1.5=6 so you know the answer is correct.

OpenStudy (anonymous):

Here is a similar question as an example, so you can see how to do the work. Mike wants to make meatloaf. His recipe uses a total of 3 pounds of meat. If he uses a 3 to 1 ratio of beef to pork, how much pork will he use? Enter your answer as a mixed number in simplest terms. 3 Beef : 1 Pork <= Convert to % by dividing by 4 75% Beef : 25% Pork <= Multiply by 3 pounds of meat desired 2.25 pounds of Beef : .75 pounds Pork